At RSM summer camp 100 students had enough food for a 30-days stay. After 2 days 40 new students joined the camp. After that 2 weeks passed and half of the students left the camp. How many additional days of food need to be ordered?
Found 2 solutions by greenestamps, ankor@dixie-net.com:
Answer by greenestamps(13241)



There is some good basic arithmetic in this problem; but the question that is asked makes no sense....

The camp starts with enough food for 100 students for 30 days.

After the first 2 days, there is enough food for 100 students for 30-2 = 28 days.

Then 40 new students joined the camp. That means the number of students increased by a factor of 140/100 = 1.4. And that means the number of days the food will last decreases by a factor of 1.4. So the food will last the 140 students for 28/1.4 = 20 days.

Then, after the next two weeks (14 days), there will be enough food for the 140 students for 20-14 = 6 days.

At that point, half the students leave the camp. That means the number of days the food still on hand will last doubles; so there is now enough food on hand for the 70 remaining students for 6*2 = 12 days.

There is nowhere in this scenario where any more food needs to be ordered....


Answer by ankor@dixie-net.com(22740)

At RSM summer camp 100 students had enough food for a 30-days stay.
100*30 = 3000 food days available
After 2 days 40 new students joined the camp.
After that 2 weeks passed and half of the students left the camp. How many additional days of food need to be ordered?
:
First 2 days: 100*2 = 200 food days consumed
Next 28 days, if no one leaves: 140 * 28 = 3920 food days required
However, 70 were not there for the last 16 days = 70 * 16 = -1120 food days
We then have 200 + 3920 - 1120 = 3000 food days, they won't have order any